Whats a comparison of the graphs: y = x and the graph of y = a(x -­ h)2+ k
a, h, and k can be any numbers.
Thanks


Found 2 solutions by ikleyn, josgarithmetic:
Answer by ikleyn(52945)

.
Check your post.

What is written there?


Answer by josgarithmetic(39632)

The only realistic comparison is they are both functions. The rest of how they relate is not very comparable. y=x is same as y=1*x, and the coefficient gives information about how y changes with x. y=a(x-h)^2+k uses the factor a to give some indication of how y changes with x.

Note that y=x has no vertex, or no special point like that. You could say, y=1(x-h)+k has a point (0,0) and gives the simplified equation y=x.